OpenStudy (anonymous):

(x/x^2−9)+(1/x+3)=(1)/(4x−12)

OpenStudy (anonymous):

Can you factor any of those expressions?

OpenStudy (jdoe0001):

hmm he hasn't said anything about factoring... he's just sharing the equation, just be happy is there =)

OpenStudy (jdoe0001):

nice equation btw

OpenStudy (anonymous):

I'll take his word that it's an equation then :p I mean that IS an equals sign

OpenStudy (anonymous):

I just have to solve for x. i think everything is already factored

OpenStudy (jdoe0001):

hmm nothing is factored btw

OpenStudy (anonymous):

i dont think you can factor anything? idk im really confused i just need to find the solution to the equation

OpenStudy (jdoe0001):

hmmm well... you'd need to brush up on your factoring it seems

OpenStudy (jdoe0001):

I'd say the same as ♂ , can you factor any of those? if not then you may want to brush it up some

OpenStudy (anonymous):

i guess x^2-9 factors but what about the rest

OpenStudy (jdoe0001):

hmmm how would you factor \(x^2-9\) anyway?

OpenStudy (anonymous):

(x+3)(x-3)

OpenStudy (jdoe0001):

hmm I see

OpenStudy (anonymous):

\[\huge \frac{x}{(x+3)(x-3)}+ \frac{1}{x+3}= \frac{1}{4x−12}\] this might help you visualize it

OpenStudy (anonymous):

like cross multiply. (x)(x-3)+........then set it equal to 1/4x-12 which it already is. then solve

OpenStudy (jdoe0001):

\(\bf \cfrac{x}{x^2-9}+\cfrac{1}{x+3}=\cfrac{1}{4x-12}\implies \cfrac{x}{(x-3)(x+3)}+\cfrac{1}{x+3}=\cfrac{1}{4x-12} \\ \quad \\ \cfrac{x+(x-3)}{(x-3)(x+3)}=\cfrac{1}{4(x-3)}\implies \cfrac{2x-3}{(x-3)(x+3)}=\cfrac{1}{4(x-3)} \\ \quad \\ 4\cancel{ (x-3) }\cdot \cfrac{2x-3}{\cancel{ (x-3) }(x+3)}=1\implies \cfrac{8x-12}{x+3}=1 \\ \quad \\ 8x-12=x+3\) and surely you can get "x" from there

OpenStudy (anonymous):

hint* fraction

OpenStudy (anonymous):

that helped a lot, thank you so much!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

yw?

OpenStudy (anonymous):

it means youre welcome

OpenStudy (anonymous):

@ambiguousmathstudent are you being sarcastic? can't tell

OpenStudy (anonymous):

yw == you're welcome

OpenStudy (anonymous):

oh. embarrassing moment for me:(

OpenStudy (anonymous):

haha dont worry bout it. thx guys